# Origin and stability of generalized Wigner crystallinity in triangular moiré systems

### v1

@@ Abstract
Generalized Wigner crystals (GWC) on triangular moiré superlattices, formed from stacking two layers of transition metal chalcogenides, have been recently observed at multiple fractional fillings of the moiré unit cell [Y. Xu et al., Nature 587, 214 - 218 (2020), X. Huang et al. Nat. Phys. 17, 715 - 719 (2021), H. Li et al., Nature 597, 650 - 654 (2021)]. Motivated by these experiments, tied with the need for an accurate microscopic description of these materials, we explore the theoretical origins of generalized Wigner crystallinity at $n=1/3$ and $2/3$ filling. We demonstrate the limitations of theoretical descriptions of these moiré GWCs relying on truncated (finite-range) electron-electron interactions instead of a long-range Coulomb interaction. We validate our findings by studying both classical and quantum mechanical effects at zero and finite temperatures. More generally, we discuss the role of charge frustration in the theoretical extended Hubbard-model phase diagram, identifying a “pinball” phase, a partially quantum melted generalized Wigner crystal with coexisting solid and liquid-like features, with no classical analog. Quantum effects also explain the small, but experimentally detectable, asymmetry in the transition temperatures of the $n=1/3$ and $2/3$ crystals. Our calculations reveal that the charge ordering temperature and the magnetic crossover temperature predicted using the long-range Coulomb interaction are correctly captured with appropriately renormalized nearest-neighbor interactions. The effective nearest-neighbor interaction strength is significantly weaker than previously reported, placing it close to a metal-insulator phase boundary. This observation has implications for future experiments that we discuss. We conclude by studying the dependence of melting temperatures on gate-to-sample separation and we also predict temperature scales at which magnetic crossovers should be observed.
@@
Introduction – The study of moiré systems has opened up multiple new avenues of research in condensed matter physics [1, 2, 3]. Previously conjectured phases of electronic matter that were attributed to just the imagination of creative theorists are now being realized and observed in these platforms, e.g., Refs. [4, 5, 6, 7, 8, 9]. The unique properties of moiré systems that allow for these proposed phases to be readily realized rely on the significant reduction in overall energy scales enabled by the moiré superlattice (while keeping strong correlations intact) combined with the tunability of carrier densities by the electrostatic gating characteristic of 2D materials. An excellent example of such theoretical and experimental synergy is shown by recent works that demonstrate how triangular lattice Hubbard model qualitatively explains the properties of moiré transition metal dichalcogenide systems (TMD) [6, 6, 10, 7, 11, 12, 13, 14, 15, 16] that display itinerant ferromagnetism (above half filling) [17, 18, 19] and antiferromagnetism from kinetic frustration (below half filling) [20, 21], with observable effects at “intermediate” temperatures [19, 22, 23].
Figure 1: (a) A schematic of $t-V$ model from Eq.(1). (b) Classical phase diagram of charge ground state configuration in the $V_{1}-V_{2}-V_{3}$ model. The $+$ indicates the presence of degenerate configurations in addition to the one mentioned. (c) Real space configuration of triangular, dimer, trimer/stripes, fourmer, and their corresponding total energy. (d, e) The energy difference of the dimer and trimer configuration compared with triangular configuration as a function of interaction cutoff $r_{c}/a$ and gate distance $d/a$, respectively. The computation is carried out on a $72\times 72$ system.
@@
While the Hubbard model description of the moiré TMD systems has undoubtedly been very successful [24, 25], a crucial aspect missing is the role the long-range nature of the Coulomb interaction can play controlling the electronic phase diagram (e.g. see the discussion on the importance of non-local/finite range interaction terms in Hubbard model in Ref. [26, 27]). In this work, we focus on the role of long-range Coulomb interaction in these systems and analyze the consequences of projecting it to a finite-range Hubbard model. Specifically, we systematically explore, at zero and finite temperature, both finite-range (FR) and long-range (LR) Hubbard models given by,
$H=-\sum_{i<j,\sigma}t_{ij}c^{\dagger}_{i,\sigma}c_{j,\sigma}+\textrm{h.c.}+U\sum_{i}n_{i,\uparrow}n_{i,\downarrow}+\sum_{i<j}V_{ij}n_{i}n_{j}$ (1)
@@
where $i,j$ refer to the sites of a (moiré) triangular lattice, $\sigma$ is the spin index, and $t_{ij}$, $U$ and $V_{ij}$ refer to the hopping, on-site interaction strength and density-density interactions respectively, see Fig.1a. Here $n_{i,\sigma}\equiv c^{\dagger}_{i,\sigma}c_{i,\sigma}$ and $n_{i}\equiv n_{i,\uparrow}+n_{i,\downarrow}$ are the usual number operators. We focus specifically on the fillings ($n$) of $1/3$ and $2/3$ to make the physics of a long-range Coulomb interaction truncation the center of our discussion, leaving an exhaustive study of other fillings for future work.
@@
The experimental platform that realizes the relevant physics for this manuscript is the moiré heterobilayer $\mathrm{WSe}_{2}$/$\mathrm{WS}_{2}$ system [6, 10, 7, 11, 12], which hosts generalized Wigner crystals (GWC) at fillings commensurate with the moiré superlattice enabled by the $\mathrm{WSe}_{2}$,$\mathrm{WS}_{2}$ lattice mismatch [14, 28]. Interest in GWCs has been fueled by recent experiments that used optical probes [6] and scanning tunneling microscopy (STM) [7] to confirm the presence of GWCs in moiré TMDs. In the STM work in particular, a state consistent with an expected $\sqrt{3}\times\sqrt{3}$ order was found in small local patches, however, vacancies in the underlying charge order with motifs that visually resemble other orders are also seen. The nature of these insulating GWC orders has been studied theoretically through a mixture of theoretical techniques starting from continuum-model or Hubbard-model-like description and employing field theory, Hartree-Fock, DMFT, Monte-Carlo, density matrix renormalization group (DMRG), or exact diagonalization (ED) techniques[29, 30, 31, 32, 33, 34, 35, 15, 36, 37].
@@
Our work addresses currently unresolved aspects of these experiments, contributing to the theoretical understanding of the moiré TMD systems and presenting both zero-temperature and finite-temperature studies in a unified framework. Our paper is structured in three parts, each addressing one precise objective. The first objective is to demonstrate that while finite-range Hubbard models have the essential elements to describe a GWC, they are generally inadequate for a quantitative explanation of the experiments. This analysis involves a general exploration of the classical ($t=0$, $U\rightarrow\infty$) and quantum mechanical phase diagram of Hamiltonian in Eq. (1) at zero temperature. The next objective is to understand the finite temperature melting of these GWCs. We build on the classical treatment that accurately modeled many aspects of the experiments [6], but did not address the role of quantum fluctuations - after all, it is not a priori clear if the melting of the GWC (at $\sim 37$ K) has any quantum origin given that the kinetic energy is $t\sim 20$ K. Our numerical calculations explain the small but experimentally detectable [6] (few Kelvin) difference in the transition temperatures of the $n=1/3$ and $2/3$ GWC, that are exactly dual to one another at the classical level. A natural outcome of the first two objectives is the partial reconciliation of conflicting parameter sets chosen by different groups [6, 13, 38]. Our third objective is to predict the extent of the stability of GWC with screening gate distance (see also Ref. [15]) - a key component of the moiré TMD experimental setup. This objective aims to pave the way for informed experiment design by determining at what temperatures the magnetic correlations present in GWC [39, 13, 40, 16, 41, 42], are most prominent and thus most easily detectable.

### v2

@@ ABSTRACT
Generalized Wigner crystals (GWC) on triangular moiré superlattices, formed from stacking two layers of transition metal chalcogenides, have been observed at multiple fractional fillings [Nature 587, 214 - 218 (2020), Nat. Phys. 17, 715 - 719 (2021), Nature 597, 650 - 654 (2021)]. Motivated by these experiments, tied with the need for accurate microscopic descriptions of these materials, we explore the origins of GWC at $n=1/3$ and $2/3$ filling. We demonstrate the general limitations of theoretical descriptions relying on finite-range, versus long-range interactions, however, we clarify why some properties are captured by an effective nearest-neighbor model. We study both classical and quantum effects at zero and finite temperatures, discussing the role of charge frustration, identifying a “pinball” phase, a partially quantum melted GWC, with no classical analog. Our work addresses several experimental observations and makes predictions for how many of the theoretical findings can be potentially realized in future experiments.
INTRODUCTION
@@
The study of moiré systems has opened up multiple new avenues of research in condensed matter physics [1, 2, 3]. Previously conjectured phases of electronic matter that were attributed to just the imagination of creative theorists are now being realized and observed in these platforms, e.g., see refs. [4, 5, 6, 7, 8, 9]. The unique properties of moiré systems that allow for these proposed phases to be readily realized rely on the significant reduction in overall energy scales enabled by the moiré superlattice (while keeping strong correlations intact) combined with the tunability of carrier densities by the electrostatic gating characteristic of 2D materials. An excellent example of such theoretical and experimental synergy is shown by recent works that demonstrate how the triangular lattice Hubbard model qualitatively explains the properties of moiré transition metal dichalcogenide systems (TMD) [6, 6, 10, 7, 11, 12, 13, 14, 15, 16] that display itinerant ferromagnetism (above half filling) [17, 18, 19] and antiferromagnetism from kinetic frustration (below half filling) [20, 21], with observable effects at “intermediate” temperatures [19, 22, 23].
Figure 1: Hamiltonian, classical phases, and energetics of the extended Hubbard model on the triangular lattice. (a) A schematic of the extended Hubbard model as in Eq. (1), with nearest neighbor hopping $t$, on-site Hubbard interaction $U$, and nearest, next-nearest and next-next-nearest neighbor interactions denoted by $V_{1},V_{2},V_{3}$ respectively. (b) Locations of lowest energy states in phase space, restricted to $n=1/3$ charge configurations shown in panel c, in the $V_{1}-V_{2}-V_{3}$ model ($V_{1}>0,U\to\infty$, $t=0$).. (c) Charge configurations for $n=1/3$ showing the triangular, dimer, trimer, fourmer crystals, from left to right, and expressions for the corresponding total energy for NN sites. (d, e) The individual energy differences of the dimer and trimer configurations, each with respect to the triangular configuration, as a function of interaction cutoff $r_{c}/a$ for a fixed $d/a=10$ (panel d), and gate distance $d/a$ for $r_{c}\rightarrow\infty$ (panel e). The calculation is carried out on a $N=72\times 72$ system with periodic boundary conditions. The y-axis label for panels d,e is the same.
@@
While the Hubbard model description of the moiré TMD systems has undoubtedly been very successful [24, 14], a crucial aspect missing is the role the long-range nature of the Coulomb interaction can play controlling the electronic phase diagram (e.g. see the discussion on the importance of non-local/finite range interaction terms in Hubbard model in refs. [25, 26, 27]). In this work, we focus on the role of long-range Coulomb interaction in these systems and analyze the consequences of projecting it to a finite-range extended Hubbard model. Specifically, we systematically explore, at zero and finite temperature, both finite-range (FR) and long-range (LR) Hubbard models given by,
$H=-\sum_{i<j,\sigma}t_{ij}c^{\dagger}_{i,\sigma}c_{j,\sigma}+\textrm{h.c.}+U\sum_{i}n_{i,\uparrow}n_{i,\downarrow}+\sum_{i<j}V_{ij}n_{i}n_{j},$ (1)
@@
where $i,j$ refer to the sites of a (moiré) triangular lattice, $\sigma$ is the spin index, and $t_{ij}$, $U$ and $V_{ij}$ refer to the hopping, on-site Hubbard interaction strength and density-density interactions respectively, see Fig.1a. $c^{\dagger}_{i,\sigma}$ and $c_{i,\sigma}$ are fermion creation and annihilation operators respectively and $n_{i,\sigma}\equiv c^{\dagger}_{i,\sigma}c_{i,\sigma}$ and $n_{i}\equiv n_{i,\uparrow}+n_{i,\downarrow}$ are number operators. We focus specifically on the fillings ($n$) of $1/3$ and $2/3$ to make the physics of the LR Coulomb interaction truncation the center of our discussion, leaving an exhaustive study of other fillings for future work.
@@
The experimental platform that realizes the relevant physics for this manuscript is the moiré heterobilayer $\mathrm{WSe}_{2}$$\mathrm{WS}_{2}$ system [6, 10, 7, 11, 12], which hosts generalized Wigner crystals (GWC) at fillings commensurate with the moiré superlattice enabled by the $\mathrm{WSe}_{2}$$\mathrm{WS}_{2}$ lattice mismatch [14, 28]. Interest in GWCs has been fueled by recent experiments that used optical probes [6] and scanning tunneling microscopy (STM) [7] to confirm their presence in moiré TMDs. In the STM work in particular, a state consistent with an expected $\sqrt{3}\times\sqrt{3}$ triangular charge order was found in small local patches, however, vacancies in the underlying charge order with motifs that visually resemble other orders are also seen. The nature of these insulating GWC orders has been studied theoretically through a combination of theoretical techniques starting from a continuum-model or Hubbard-model-like description and employing field theory, Hartree-Fock, dynamical mean field theory, Monte-Carlo, density matrix renormalization group (DMRG), and exact diagonalization (ED) [26, 29, 30, 31, 32, 33, 15, 34, 35]. We note that the term GWC is often used interchangeably with charge density wave (CDW), however, the difference is that CDW emerges as an instability due to Fermi surface nesting whereas in a GWC the charge localization is interaction driven and the kinetic energy is strongly quenched.
@@
Our work addresses currently unresolved aspects of these experiments, contributing to the theoretical understanding of the moiré TMD systems and presenting both zero-temperature and finite-temperature studies in a unified framework. Our paper is structured in three parts, each addressing one precise objective. The first objective is to demonstrate that while FR Hubbard models have the essential elements to describe a GWC, they are typically inadequate for a quantitative explanation of the experiments if the LR tail is rigidly truncated. This analysis involves a general exploration of energetically competitive phases at the classical ($t=0$, $U\rightarrow\infty$) and quantum mechanical level for the Hamiltonian in Eq. (1) at zero temperature. The next objective is to understand the finite temperature melting of these GWCs. We build on the classical treatment that accurately modeled many aspects of the experiments [6], but did not address the role of quantum effects - after all, it is not a priori clear if the melting of the GWC (at $\sim 37$ K) has any quantum origin given that the kinetic energy is $t\sim 20$ K. Our numerical calculations explain the small but experimentally detectable [6] (few Kelvin) difference in the transition temperatures of the $n=1/3$ and $2/3$ GWC, that are exactly dual to one another at the classical level. A natural outcome of the first two objectives is the partial reconciliation of conflicting parameter sets chosen by different groups [6, 13, 36]. Our third objective is to predict the extent of the stability of GWC with screening gate distance (see also ref. [15]) - a key component of the moiré TMD experimental setup. This objective aims to pave the way for informed experiment design by determining at what temperatures the magnetic correlations present in GWC [37, 13, 38, 16, 39, 40], are most prominent and thus most easily detectable.
RESULTS
Classical phase diagram and competing charge-ordered states